# ppt: A Process-based Preference Learning Framework for Self Improving Table Question Answering Models

### v1

@@ 5 Results and Discussions
Figure 3 shows the Exact Match of different methods on in-domain datasets, averaged across models. Per model results can be found in Appendix A.7.
Figure 3: Comparing our ppt with baselines using Exact Match (EM). Results are averaged across models. RFT and FDPO stand for rejected sampling fine-tuning and full chain DPO. We experiment with value functions: mc-b and mix. Dashed lines show performances of fine-tuned TQA models $M_{ft}$ before applying self-improvement methods.
ppt effectively improves performance of TQA models.
As Figure 3 shows, applying ppt enhances the performance of $M_{ft}$ by 3.5% on average on in-domain datasets. The gains are more obvious on WTQ (5%) compared to TabFact (2%). This might be attributed to dataset features: TabFact is a binary classification dataset, thus it is easier for models to achieve high performance and harder to be further improved. When evaluating on out-of-domain datasets, we witness an average of 2.2% performance gain after applying our framework ($\uparrow$ 2.2% for WikiSQL, $\uparrow$ 2.4% for SCITAB and $\uparrow$ 2.1% for CRT). The findings show the genealisibility of our method on out-of-domain data.
Process-based preference learning is effective and efficient for self-improving TQA models. Comparing ppt with RFT and FDPO, we find RFT does not improve performance. In contrast, preference learning effectively self-improves models. Moreover, process-based preference learning (ppt) outperforms coarse-grained full chain learning (FDPO) on two out of three datasets. Comparing different value functions, we find ppt with continuous MC estimation to work better than other methods. This is an important finding as it shows that simply using $M_{ft}$ to estimate state values with a high threshold $\tau=0.9$ to control pair qualities can work similarly or better than additionally querying an LLM judge. It significantly boosts the efficiency of the framework: we find using ppt takes 10 times less time than using ppt(mix) when sampling the same number of data. In a manual evaluation of step correctness (Shown in Appendix A.8), we do not find differences between ppt and ppt(mix). This suggests ppt’s effectiveness and efficiency are not compromised for reasoning correctness. We also provide an analysis of threshold impact in Appendix A.9, showing the necessity of picking a relatively high threshold for distinguishing step pairs during preference learning.
@@
Table 1: Exact Match of TQA models. $M_{ft}$ and $M_{si}$ refers to fine-tuned and self-improved TQA models respectively. T-LlaMA refers to TableLlaMA and T-SQLify refers to TabSQLify. Framework results (the last two rows) are obtained using GPT-3.5 as the backbone. SoTA TQA results are obtained from previous work Zhou et al. (2024); Zhang et al. (2023a); Wu and Feng (2024); Nahid and Rafiei (2024).
@@
Self-improved TQA models achieve competitive results compared with more complicated SoTA approaches with five times less inference time. Table 5 shows results for current TQA models. For the sake of space, we only show results from Qwen-2.5. Full results containing LlaMA-3.1 are shown in Appendix A.7. The first four rows of Table 5 compare small-size fine-tuned TQA models, and the last two rows show SoTA training-free frameworks back-boned by GPT-3.5. We find $M_{si}$, enhanced using ppt outperforms existing TQA models. More importantly, it achieves competitive performance compared with larger and more complex frameworks where tools and agentic collaboration are involved. On SCITAB, $M_{si}$ even achieves the best performance. We emphasize the inference efficiency of $M_{si}$: it requires eight times less inference time than MACT and five times less than TabSQLify, yet provides competitive results.
@@ 6 Conclusions
In this study, we provided the first self-improvement framework for TQA, using process-based preference learning. Our framework effectively improved the performance of TQA models up to 5% and 2.4% on in-domain datasets and out-of-domain datasets respectively. The resulting models demonstrated competitive performances compared to state-of-the-art TQA systems, which are dependent on large LLMs and tool usages. Yet, it requires five times less inference time.
Limitation
While our method effectively enhances the performance of small-size TQA models, its impact on large TQA models remains unexplored. To the best of our knowledge, current fine-tuned TQA models only focus on small-size LLMs. Future work can explore an efficient training strategy for large fine-tuned TQA models. Secondly, we limit the task in our study to only TQA, while there exist other table-related tasks, such as table summarization. Thirdly, although our framework supports iterative self-learning, the present work only demonstrates the effectiveness of the first iteration, leaving multi-iteration evaluations for future study. As the datasets we used in this study are originally sourced from Wikipedia, scientific paper and statistical reports, we do not observe any potential risks from the datasets.
@@ A.1 Prompts
Figure 6, 7 and 8 show prompts for generating a full reasoning trace, completing a reasoning trace and LLM judge evaluation for a reasoning trace.
@@ A.2 Datasets
Table 2 shows number of instances and domains for the test data we used.
A.3 Fine-tune TQA Models
Following previous work Wu and Feng (2024); Zhang et al. (2025b), we prompt Deepseek-V3 DeepSeek-AI et al. (2025) for reasoning chains that lead to correct answers to fine-tune our TQA models. Specially, we sample 2.4k, 1.5k and 2.3k data from the training set of WTQ Pasupat and Liang (2015), TabFact Chen et al. (2020) and HiTab Cheng et al. (2022). Then we prompt Deepseek-V3 to generate reasoning chains and answers. We only keep reasoning chains leading to correct answers for fine-tuning. This results in 1612 instances from WTQ, 1425 instances from TabFact, and 1277 from HiTab. In total, we collect 4314 instances.
A.4 Baselines
Baselines for Self-Improvement Methods.
To collect data for rejected sampling fine-tuning (RFT), we select the reasoning chains leading to the correct answer: $D_{rft}=\{(r_{w})^{i}\}_{i=1}^{|D_{rft}|}$. We fine-tune TQA models with the following loss function for self-improvement:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(s_{0},r_{w})\sim\mathcal{D}_{rft}}\left[\log\pi_{\theta}(r_{w}|s_{0})\right]$ (3)
To obtain full reasoning chain DPO data, we selecting reasoning chains leading to correct as preferred traces and those leading to incorrect answers as not preferred: $D_{fdpo}=\{(s_{0},r_{w},r_{l})^{i}\}_{i=1}^{|D_{fdpo}|}$, where $r_{w}$ and $r_{l}$ refer to the preferred and not preferred reasoning traces respectively. The loss function can be represented as:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(u,r_{w},r_{l})\sim\mathcal{D}_{dpo}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(r_{w}|u)}{\pi_{ref}(r_{w}|u)}\right)\right.\\ \left.-\beta\log\frac{\pi_{\theta}(r_{l}|u)}{\pi_{ref}(r_{l}|u)}\right]\end{aligned}$
Baselines for TQA Systems.
TableLlaMA Zhang et al. (2023a) is an end-to-end TQA system fine-tuned on 2.6M existing table-related data. It uses LlaMA-2-7B as a base model and directly outputs answers without a reasoning chain. Protrix Wu and Feng (2024) is fine-tuned with around 4k table-related data with reasoning chain generated from GPT-4. It also uses LlaMA-2-7B as the base model. MACT Zhou et al. (2025) is a training-free TQA framework, leveraging tools and agent collaboration. Specifically, a planning agent outputs step-wise plans to decompose questions, and a coding agent and tools solve each step. TabSQLify Nahid and Rafiei (2024) decomposes tables into relevant sub-tables with SQL query generation and execution. Then, sub-tables and questions are passed to LLMs to obtain final answers.
@@ A.4 Baselines
Table 2: Test data statistics.
A.5 Sampled Dataset Statistics
Table 3 shows sampling size for each method. We find mc-b results in the most data while RFT the least.
@@ A.6 Hyper-parameters
A.7 Additional Results
Table 5 shows different models’ performance on the six investigated datasets.
@@ A.7 Additional Results
A.8 Reasoning Chains Analysis
We sample 100 reasoning chains leading to correct answers generated from models using ppt and models using mix as the value function. We manually examine the correctness of the reasoning chains. Among the 100 instances, we exclude 8 instances where either the answers are incorrect or the questions are ambiguous. We find similar accuracies of the reasoning chains generated from the aforementioned methods, with 95.7 and 94.6, respectively. This suggests the two methods do not differ much in terms of leading to correct reasoning chains. Nevertheless, wrong reasoning chains leading to correct answers still exist, possibly due to overly complex table inputs. An error case is shown in Figure 9.
WTQTabFactHiTabWikiSQLSCITABCRT5050505055555555606060606565656570707070757575758080808085858585DatasetsEMThresholds Comparison with Soft EstimationSoft Estimation ($\tau$=0.5)Soft Estimation ($\tau$=0.7)Soft Estimation ($\tau$=0.9)
@@ A.8 Reasoning Chains Analysis
Figure 5: Confusion matrix between LLM judge evaluation and Monte Carlo estimation with binary labels.
A.9 Threshold Analysis
@@
A.10 Dataset Licenses
WTQ Pasupat and Liang (2015), HiTab Cheng et al. (2022) and WikiSQL Zhong et al. (2017) are under the license of CC-BY-SA-4.0111https://creativecommons.org/licenses/by-sa/4.0/, BSD-3 Clause [^2] https://opensource.org/license/bsd-3-clause and C-UDA [^3] https://github.com/microsoft/HiTab?tab=License-1-ov-file respectively. TabFact Chen et al. (2020), CRT Zhang et al. (2023b) and SCITAB Lu et al. (2023) are under the MIT [^4] https://opensource.org/license/mit license.
@@ A.10 Dataset Licenses
Figure 9: Wrong reasoning chain generated by a self-improved model (Qwen-2.5-7B) using ppt. The first wrong step is highlighted with red.

### v2

@@ 4 Results and Discussions
Figure 3 shows the Exact Match of different methods on in-domain datasets, averaged across models. Per-model results can be found in Appendix A.5.
Figure 3: Comparing p2-TQA with baselines using Exact Match. Results are averaged across models. RFT and FDPO stand for rejected sampling fine-tuning and full-chain DPO, respectively. We experiment with several value functions: Self-Exp (Self-Exploration), mc-b (Monte Carlo with binary values), and mix (a combination of LLM-as-a-judge and mc-b). Dashed lines show performances of fine-tuned TQA models $M_{ft}$ before applying self-improvement methods.
p2-TQA effectively improves the performance of TQA models.
As Figure 3 shows, applying p2-TQA enhances the performance of $M_{ft}$ by 3.5% on average on in-domain datasets. The gains are more obvious on WTQ (5%) compared to TabFact (2%). This might be attributed to dataset features: TabFact is a binary classification dataset, thus it is easier for models to achieve high performance and harder to further reach improvements. When evaluating on out-of-domain datasets, we witness an average of 2.2% performance gain after applying our framework ($\uparrow$ 2.2% for WikiSQL, $\uparrow$ 2.4% for SCITAB, and $\uparrow$ 2.1% for CRT). These findings demonstrate the generalisability of our method on out-of-domain data. The performance gains on both in-domain and out-of-domain datasets are significant and can be observed with inference sampling over multiple runs. Detailed results are shown in Appendix A.5.
@@
Table 1: Exact Match of TQA models. $M_{si}$ refers to self-improved models. T-LlaMA and T-SQLify refer to TableLlaMA and TabSQLify, respectively. Framework results (the last two rows) are obtained using GPT-3.5 as the backbone. State-of-the-art TQA results are obtained from previous work zhou-etal-2024-freb; Zhang2023TableLlamaTO; wu-feng-2024-protrix; nahid-rafiei-2024-tabsqlify.
p2-TQA delivers competitive results against baselines, highlighting the effectiveness of pairing a lightweight value function with stringent filtering for self-improving models.
Comparing p2-TQA with RFT and FDPO, we find that our framework leads to higher improvements. The effect is more obvious on in-domain datasets than out-of-domain datasets, as results in Appendix A.5 show. Though FDPO is generally computationally cheaper than p2-TQA, i.e., under the same token budgets, it generates more training instances. We find that the performance of p2-TQA improves when training example sizes increase. In contrast, more examples do not necessarily lead to better performance using FDPO, suggesting early saturation. More analysis is presented in Appendix A.6.
When comparing against methods using different value functions, we observe that p2-TQA greatly outperforms Self-Explore. It also shows advantages over methods that simply use a binary value function (mc-b), or combining it with an LLM judge (mix). Notably, p2-TQA takes 10 times less time than using mix when sampling the same amount of data. This demonstrates that our method delivers strong performance while being efficient. Interestingly, the efficiency is not compromised for reasoning correctness. We evaluate models fine-tuned using data generated by p2-TQA and mix with regard to step correctness and find that the two methods achieve similar accuracy (95.7% vs. 94.6%). Detailed analysis can be found in Appendix A.7. We provide an analysis of threshold impact in Appendix A.8, showing the necessity of picking a relatively high threshold for data filtering.
Self-improved TQA models achieve competitive results compared to complicated state-of-the-art approaches, with five times less inference time.
@@
Table 1 shows results for current TQA models. first four rows of Table 1 compare small-size fine-tuned TQA models, and the last two rows show state-of-the-art training-free frameworks back-boned by GPT-3.5. We find that both the Qwen and LlaMA models, enhanced using p2-TQA, outperform existing TQA models. More importantly, both self-improved models achieve competitive performance compared to larger and more complex frameworks where tools and agentic collaboration are involved. On SCITAB, $M_{si}$ with the Qwen even achieves the best performance. Apart from the competitive task performance, we emphasize the inference efficiency of self-improved models: they require eight times less inference time than MACT and five times less than TabSQLify.
Model Retrieval Reasoning Total
# instances 1133 451 1584
Qwen ($M_{ft}$) 71.32 36.36 61.36
Qwen ($M_{ft}$) +p2-TQA 78.02 41.24 67.55
LlaMA ($M_{ft}$) 79.44 45.23 69.70
LlaMA ($M_{ft}$) +p2-TQA 81.38 42.57 70.32
Table 2: Exact Match of models with and without p2-TQA, evaluated across different question types on the HiTab test set.
Applying p2-TQA generally improves accuracy across question types, table sizes, and step correctness.
We compare models fine-tuned with p2-TQA against those without, from two perspectives: We take a closer look at question type and table size, inspired by zhou-etal-2025-texts. For question type analysis, we categorize questions into those requiring only retrieval and those requiring reasoning in addition. We evaluate models on the HiTab dataset, which provides explicit question type annotations. As shown in Table 2, the enhanced Qwen-2.5-7B model outperforms its pre-trained counterpart in both categories, whereas the LLaMA-3.1-8B model shows notable gains primarily in retrieval questions. For table size analysis, we partition tables into three bins based on token count and compute EM accuracy for each bin. Figure 4 reports results averaged over in-domain datasets, revealing that self-improved models consistently achieve higher accuracy across all table sizes.
Finally, we examine step correctness by comparing models with ($M_{si}$) and without ($M_{ft}$) p2-TQA. We randomly sample 50 instances from HiTab and manually examine the correctness of reasoning steps for both Qwen and LLaMA, yielding a total of 200 reasoning chains (50 instances × 2 models × 2 variants) and 245 steps for $M_{ft}$ versus 239 steps for $M_{si}$. Models enhanced with p2-TQA demonstrate higher step accuracy than those without (83% vs. 75%), with most improvements arising from reduced errors in planning and numerical reasoning.
Figure 4: Exact Match of models with and without p2-TQA, evaluated across different table sizes and averaged over in-domain datasets. Qw and LM show the performance of Qwen MftM{ft} and LlaMA MftM{ft} respectively. Instances are grouped into three bins by table token count: <500<500, 50050010001000, and $\geq 1000$.
@@ 5 Conclusions
We have introduced a self-improvement framework p2-TQA that uses process-based preference learning. Our framework effectively improves the performance of TQA models by up to 5%. The resulting models demonstrate competitive performance compared to state-of-the-art TQA systems, which depend on huge LLMs and tool usage. Yet, models enhanced with p2-TQA require five times less inference time.
Limitations
First, While our method effectively enhances the performance of small-sized TQA models, its impact on large TQA models remains unexplored. To the best of our knowledge, current fine-tuned TQA models only focus on small-sized LLMs. Future work can explore an efficient training strategy for large fine-tuned TQA models. Second, we limit the task in our study to only TQA, while there exist other table-related tasks, such as table summarization. Third, although our framework supports iterative self-learning, the present work only demonstrates the effectiveness of the first iteration, leaving multi-iteration evaluations for future study. As the datasets we used in this study are originally sourced from Wikipedia, scientific papers, and statistical reports, we do not observe any potential risks from the datasets.
@@ A.1 Prompts
Figure 6, 7, and 8 show prompts for generating a full reasoning trace, completing a reasoning trace, and LLM judge evaluation for a reasoning trace.
@@ A.2 Datasets
Table 3 shows the number of instances and domains for the test data we used. WTQ pasupat-liang-2015-compositional, HiTab cheng-etal-2022-hitab and WikiSQL zhongSeq2SQL2017 are under the license of CC-BY-SA-4.0222https://creativecommons.org/licenses/by-sa/4.0/, BSD-3 Clause [^3] https://opensource.org/license/bsd-3-clause and C-UDA [^4] https://github.com/microsoft/HiTab?tab=License-1-ov-file respectively. TabFact 2019TabFactA, CRT zhang-etal-2023-crt and SCITAB lu-etal-2023-scitab are under the MIT [^5] https://opensource.org/license/mit license.
@@
Table 3: Test data statistics. The second column shows the number of test instances in each dataset.
A.3 Sampled Dataset Statistics
Table 4 shows the sampling size for each method. We find mc-b results in the most data while RFT the least.
@@ A.4 Hyper-parameters
A.5 Additional Results
Table 6 shows different models’ performance on the six investigated datasets under greedy decoding. To validate the effectiveness of applying p2-TQA, we set the temperature to 0.8 during inference. We report the mean and standard deviation over five runs in Table 7.
@@ A.5 Additional Results
Models WTQ TabFact HiTab WiKiSQL SCITAB CRT
Qwen-2.5-7B ($M_{ft}$) 57.00 $\pm{0.46}$ 81.85 $\pm{0.22}$ 61.76 $\pm{0.22}$ 67.68 $\pm{0.22}$ 52.16 $\pm{1.13}$ 46.92 $\pm{1.15}$
    +RFT 51.08 $\pm{0.60}$ 80.28 $\pm{0.32}$ 56.92 $\pm{0.55}$ 61.77 $\pm{0.08}$ 49.90 $\pm{0.83}$ 46.02 $\pm{0.60}$
    +FDPO 58.54 $\pm{0.23}$ 84.32 $\pm{0.28}$ 61.40 $\pm{0.49}$ 68.20 $\pm{0.23}$ 56.98 $\pm{0.80}$ 49.64 $\pm{1.40}$
    mc-b 57.36 $\pm{0.33}$ 82.74 $\pm{0.04}$ 62.63 $\pm{0.65}$ 66.39 $\pm{0.17}$ 52.71 $\pm{0.71}$ 48.71 $\pm{1.00}$
    mix 60.43 $\pm{0.57}$ 84.79 $\pm{0.10}$ 62.58 $\pm{0.41}$ 68.59 $\pm{0.14}$ 51.98 $\pm{1.05}$ 52.09 $\pm{0.37}$
    p2-TQA 60.42 $\pm{0.29}$ 84.81 $\pm{0.21}$ 65.05 $\pm{0.80}$ 68.90 $\pm{0.13}$ 55.08 $\pm{1.04}$ 50.99 $\pm{0.80}$
LlaMA-3.1-8B ($M_{ft}$) 57.67 $\pm{0.38}$ 82.82 $\pm{0.29}$ 63.18 $\pm{0.80}$ 62.17 $\pm{0.18}$ 49.98 $\pm{1.34}$ 47.86 $\pm{0.65}$
    +RFT 58.27 $\pm{0.47}$ 81.44 $\pm{0.15}$ 63.14 $\pm{0.70}$ 65.47 $\pm{0.22}$ 47.83 $\pm{0.98}$ 47.91 $\pm{1.61}$
    +FDPO 62.18 $\pm{0.41}$ 85.52 $\pm{0.34}$ 65.83 $\pm{0.43}$ 68.12 $\pm{0.18}$ 52.94 $\pm{0.39}$ 50.41 $\pm{0.60}$
    mc-b 60.92 $\pm{0.18}$ 83.98 $\pm{0.17}$ 65.86 $\pm{0.69}$ 65.86 $\pm{0.30}$ 52.25 $\pm{1.54}$ 49.07 $\pm{0.83}$
    mix 59.57 $\pm{0.62}$ 84.99 $\pm{0.07}$ 62.11 $\pm{0.44}$ 63.23 $\pm{0.19}$ 51.42 $\pm{0.63}$ 50.05 $\pm{1.09}$
    p2-TQA 62.66 $\pm{0.55}$ 85.71 $\pm{0.23}$ 66.49 $\pm{0.79}$ 65.83 $\pm{0.26}$ 53.76 $\pm{0.56}$ 49.15 $\pm{1.20}$
Table 7: Exact Match accuracies of models fine-tuned with different strategies and value functions, generated with sampling. p2-TQA significantly improve fine-tuned model. Compared to baselines, it achieves competitive performance across in-domain datasets.
Method 2k 4k 6k 8k 12k
[…]
Table 8: Exact Match against varying training sizes. Results are obtained by averaging across three runs and three in-domain datasets using Qwen-2.5-7B.
A.6 Cost Effectiveness Analysis
Note that step-wise sampling requires higher computing budgets than full chain sampling; we conduct a cost-effectiveness analysis over FDPO and p2-TQA. We do that by examining models’ performance under different computing budgets (approximately by training instances). We first calculate the average number of tokens needed to generate an instance for different methods. This results in approximately 3k for FDPO and 10K for p2-TQA. We fine-tune models with varying sizes of training samples. This not only allows us to compare FDPO and p2-TQA under the same computing budgets, but also demonstrates each method’s sensitivity to training size. Table 8 shows the results. We observe that for FDPO, adding more training data does not necessarily improving models’ performance (69.1–>67.1). In contrast, scaling training sizes remains effective for p2-TQA. This suggests that though FDPO generates more training instances than our method under the same budgets, the real effect of the generated data on performance is limited.
A.7 Reasoning Chains Analysis
We sample 100 reasoning chains leading to correct answers generated from models using p2-TQA and models using mix as the value function. We manually examine the correctness of the reasoning chains. Among the 100 instances, we exclude 8 instances where either the answers are incorrect or the questions are ambiguous. We find similar accuracies of the reasoning chains generated from the aforementioned methods, with 95.7 and 94.6, respectively. This suggests the two methods do not differ much in terms of leading to correct reasoning chains. Nevertheless, wrong reasoning chains leading to correct answers still exist, possibly due to overly complex table inputs. An error case is shown in Figure 9.
WTQTabFactHiTabWikiSQLSCITABCRT50505555606065657070757580808585DatasetsEMThresholds Comparison with Soft EstimationSoft Estimation ($\tau$=0.5)Soft Estimation ($\tau$=0.7)Soft Estimation ($\tau$=0.9)
@@ A.7 Reasoning Chains Analysis
A.8 Threshold Analysis
@@
Figure 9: Wrong reasoning chain generated by a self-improved model (Qwen-2.5-7B) using p2-TQA. The first wrong step is highlighted with red.